### v1

# MMhops-R1: Multimodal Multi-hop Reasoning

Abstract
The ability to perform multi-modal multi-hop reasoning by iteratively integrating information across various modalities and external knowledge is critical for addressing complex real-world challenges. However, existing Multi-modal Large Language Models (MLLMs) are predominantly limited to single-step reasoning, as existing benchmarks lack the complexity needed to evaluate and drive multi-hop abilities. To bridge this gap, we introduce MMhops, a novel, large-scale benchmark designed to systematically evaluate and foster multi-modal multi-hop reasoning. MMhops dataset comprises two challenging task formats, Bridging and Comparison, which necessitate that models dynamically construct complex reasoning chains by integrating external knowledge. To tackle the challenges posed by MMhops, we propose MMhops-R1, a novel multi-modal Retrieval-Augmented Generation (mRAG) framework for dynamic reasoning. Our framework utilizes reinforcement learning to optimize the model for autonomously planning reasoning paths, formulating targeted queries, and synthesizing multi-level information. Comprehensive experiments demonstrate that MMhops-R1 significantly outperforms strong baselines on MMhops, highlighting that dynamic planning and multi-modal knowledge integration are crucial for complex reasoning. Moreover, MMhops-R1 demonstrates strong generalization to tasks requiring fixed-hop reasoning, underscoring the robustness of our dynamic planning approach.
Codehttps://github.com/taoszhang/MMhops-R1
Introduction
Figure 1: Comparison of reasoning types. (A) Previous KB-VQA: Single-step visual recognition followed by knowledge retrieval. (B1) Bridging reasoning: Multi-step sequential inference on a single image. (B2) Comparison reasoning: Cross-image entity identification and comparative analysis.
With continuous advancement in reasoning capabilities, Large Language Models (LLMs) like OpenAI’s o1 (Jaech et al. 2024), DeepSeek-R1 (Guo et al. 2025), and Kimi-k1.5 (Team et al. 2025) demonstrate strong performance in complex problem-solving by extending chain-of-thought reasoning during inference. Multimodal large language models (MLLMs), by inheriting the reasoning abilities or adopting similar training paradigms, achieve significant progress in integrating visual understanding and language reasoning (Xu et al. 2024; Peng et al. 2025; Zhang et al. 2025; Zheng et al. 2025). However, current multimodal reasoning research primarily focuses on stimulating intrinsic model capabilities, such as spatial reasoning (Zhou et al. 2025), object detection (Chen et al. 2025a; Liu et al. 2025), and mathematical reasoning (Meng et al. 2025; Leng et al. 2025). By contrast, complex real-world problems typically require the integration of multimodal reasoning with external knowledge retrieval through multi-turn interactions, enabling multimodal multi-hop reasoning. For example, Figure 1(B1) requires the model to extract information from the image, retrieve relevant external knowledge, and perform multi-step reasoning to reach the answer. Figure 1(B2) requires identifying entities across multiple images, retrieving corresponding knowledge, and conducting quantitative reasoning.
Despite progress, existing Visual Question Answering (VQA) datasets remain insufficient for multimodal multi-hop reasoning. Current datasets are limited in both visual and textual reasoning depth: standard VQA datasets typically require only single-step visual understanding (Goyal et al. 2017; Hudson and Manning 2019; Singh et al. 2019), while several knowledge-based VQA datasets (Marino et al. 2019; Schwenk et al. 2022; Lerner et al. 2022; Chen et al. 2023) introduce external knowledge retrieval to increase complexity. However, as shown in Figure 1(A), models usually use one step of visual recognition and one step of text retrieval to answer, without constructing complex reasoning chains. E-VQA (Mensink et al. 2023) extends questions to two-hop reasoning, but this extension remains restricted to the textual domain and features a fixed reasoning path length, lacking multimodal integration and diverse reasoning types. These limitations make existing datasets inadequate for effectively supporting model training and evaluation in complex multimodal multi-hop reasoning tasks.
Based on these challenges, we propose MMhops, a novel large-scale Multimodal Multi-hop reasoning dataset that systematically increases reasoning depth in both visual and textual dimensions. MMhops features two types of reasoning tasks: Bridging reasoning and Comparison reasoning. Bridging reasoning starts from a single image and requires the model to perform multi-step chain reasoning, with each step building on the previous one, supporting reasoning depths from two hops and beyond. Comparative reasoning is based on multiple images, requiring the model to identify multiple visual entities and compare their shared attributes, involving cross-image information integration and comparative analysis. Both task types demand deep reasoning abilities in visual understanding and textual inference, enabling the model to decompose complex questions and dynamically construct answers through multi-round interactions, thus providing a comprehensive evaluation of multimodal reasoning and knowledge integration capabilities.
To address the challenges of multimodal multi-hop reasoning, we propose MMhops-R1, the first framework to leverage reinforcement learning (RL) for multimodal multi-hop reasoning. MMhops-R1 adopts a dynamic interaction strategy that overcomes the limitations of fixed processes in conventional multimodal Retrieval-Augmented Generation (mRAG) frameworks. Specifically, the model supports three core actions: 1) selecting an input image and invoking the image retriever; 2) submitting a text query to the text retriever; and 3) generating answer based on the current information. With a tailored reward mechanism, MMhops-R1 can autonomously select reasoning strategies, dynamically adjust reasoning depth according to question complexity, and adaptively plan the reasoning path.
We evaluate MMhops-R1 on the proposed MMhops benchmark against four categories of strong baselines: open-source MLLMs, multi-hop RAG, multimodal RAG, and proprietary MLLMs. Results demonstrate the profound effectiveness of our proposed RL-driven framework for dynamic mRAG and underscores two critical requirements for complex multi-modal reasoning: the ability to integrate multimodal external knowledge and to dynamically interact with a retrieval system. Furthermore, MMhops-R1 shows strong generalization, achieving robust performance on the single-hop questions from INFOSEEK and the two-hop questions from E-VQA. These findings validate our contributions and highlight the potential of our approach to drive future research in multi-modal multi-hop reasoning.
Our contributions are summarized as follows:
  • We introduce MMhops, the first large-scale benchmark for multimodal multi-hop reasoning, requiring the synthesis of diverse visual and textual information across various reasoning depths.
  • We propose MMhops-R1, a novel mRAG framework that leverages reinforcement learning to optimize the model, enabling it to dynamically interact with multiple retrievers and adaptively plan the reasoning path.
  • We set a new state-of-the-art on multimodal multi-hop reasoning tasks, demonstrating the superiority of our dynamic mRAG framework over existing methods.
Figure 2: The multi-stage construction process for the MMhops dataset.
Related Work
Knowledge-Based VQA. To advance VQA beyond perception towards more complex reasoning, the task of KB-VQA was introduced, which requires models to incorporate external knowledge. However, prominent KB-VQA datasets like OK-VQA (Marino et al. 2019) and A-OKVQA (Schwenk et al. 2022) were largely confined to commonsense knowledge or simple facts. Subsequent efforts, including ViQuAE (Lerner et al. 2022) and INFOSEEK (Chen et al. 2023), expanded the knowledge domain to large-scale corpora such as Wikipedia. Nonetheless, these datasets predominantly feature questions solvable via a two-step process: identifying a visual entity and executing a single query against a knowledge base. While the recent E-VQA (Mensink et al. 2023) dataset introduced textual two-hop reasoning, its reasoning chains are confined to the textual modality and a fixed length. In contrast, the MMhops dataset is the first to systematically require multi-hop reasoning across both visual and textual modalities, featuring diverse, variable-length reasoning paths, demanding a more profound integration of multi-modal information.
Multimodal and Multi-hop RAG. Early mRAG frameworks (Caffagni et al. 2024; Yan and Xie 2024; Zhang et al. 2024) typically employ a static, single-step pipeline: they first retrieve relevant documents based on the initial query and then feed them to the generator. A key limitation of these approaches is their reliance on a static, pre-defined process, which lacks the flexibility to adapt to queries of varying complexity. While recent work like OmniSearch (Comanici et al. 2025) introduces a planning agent, it relies on manually engineered prompts or supervised fine-tuning, which does not equip the model with the intrinsic capability to learn complex reasoning policies autonomously. In parallel, multi-hop RAG has emerged in the unimodal text domain to address similar challenges. To move beyond fixed reasoning chains, methods such as Search-R1 (Jin et al. 2025) and ReSearch (Chen et al. 2025b) leverage reinforcement learning (RL) with algorithms like GRPO (Shao et al. 2024) and PPO (Schulman et al. 2017) to train an agent that learns a dynamic retrieval policy. However, these powerful RL-based paradigms have thus far been confined to the textual modality. Our work, MMhops-R1, bridges this divide by extending this RL-based paradigm to the mRAG domain, training an agent to strategically orchestrate retrieval and reasoning across both visual and textual knowledge sources.
Dataset Scale Visual Reasoning Text Reasoning Total Reasoning Multi-image Knowledge Source
OK-VQA (Marino et al. 2019) 14K 1 1 2 Factoid
A-OKVQA (Schwenk et al. 2022) 24.9K 1 1 2 Common sense/Factoid
ViQuAE (Lerner et al. 2022) 3.7K 1 1 2 Wikipedia
INFOSEEK (Chen et al. 2023) 1.35M 1 1 2 Wikipedia
E-VQA (Mensink et al. 2023) 1M 1 1-2 2-3 Wikipedia
MMhops 31.1K 1-2 2-3 3-4 Wikipedia
Table 1: Comparison with Existing Knowledge-based VQA Datasets.
Statistic Dimension Value Percentage
Dataset Scale
Total VQA Samples 31,117 100.0%
Bridging VQA Samples 26,437 85.0%
Comparison VQA Samples 4,680 15.0%
Number of Questions Involved 20,483 -
Number of Entities Involved 8,832 -
Number of Images Involved 28,256 -
Reasoning Complexity
Requiring External Knowledge 31,117 100.0%
3 steps 22,016 70.8%
4 steps 9,101 29.2%
Content Characteristics
Average Question Length (words) 17.3 -
Average Answer Length (words) 1.6 -
Answer Type Distribution
Entity-type Answers 5,923 19.0%
Temporal Answers 5,016 16.1%
Numerical Answers 20,178 64.9%
Table 2: Statistics of MMhops Dataset
MMhops Dataset
In this section, we present MMhops, a large-scale multimodal multi-hop reasoning dataset. MMhops requires models to: (1) interact with diverse external knowledge sources for targeted retrieval; (2) perform multi-step reasoning by dynamically integrating and updating knowledge across multiple retrieval and reasoning steps; (3) align and combine information from multiple images for cross-image and cross-modal reasoning. Through a well-designed data construction and evaluation framework, MMhops serves as an important resource for multimodal reasoning research.
MMhops Construction
The MMhops dataset is built based on the Wikipedia Knowledge Base (KB), with an automated data annotation and quality filtering process designed using powerful language models like GPT-4o (Hurst et al. 2024). The dataset includes two core reasoning types: Bridging Questions and Comparison Questions, covering various reasoning depths and different numbers of image inputs.
Bridging Dataset
Bridging Questions begin with the visual information from a single image and progressively link relevant entities and knowledge through multi-step chain reasoning. We generate them iteratively by starting with single-hop questions and progressively increasing the reasoning depth, a process depicted in Figure 2 (a).
Initialization Phase:
  1. 0.
    Data Collection: Gather existing single-hop knowledge-based datasets $(V,Q^{1},A^{1})$ as the foundation for constructing multi-hop reasoning chains.
Iterative Expansion Phase:
  1. 1.
    Bridging Entity Identification: From the current n(1)n\ ($\geq$ 1) hop dataset $(V,Q^{n},A^{n})$, select samples where the answer corresponds to a Wikipedia entity $(V_{i},Q^{n}_{i},A^{n}_{i})$, excluding vague entity types such as numbers or years. The answer $A^{n}_{i}$ is designated as the bridging entity for subsequent reasoning chains.
  2. 2.
    Sub-question Generation: Using the Wikipedia page of the bridging entity $A^{n}_{i}$, prompt a large language model to generate a knowledge-based question $q^{n+1}_{i}$, ensuring that the entity $A^{n}_{i}$ is explicitly mentioned in the question and labeling the answer $A^{n+1}_{i}$.
  3. 3.
    Question Quality Control: Verify that the sub-question $q^{n+1}_{i}$ meets the criteria of an independent single-hop question, meaning that removing the entity $A^{n}_{i}$ from $q^{n+1}_{i}$ should render the question unanswerable.
  4. 4.
    Question Merging: Merge the sub-question $q^{n+1}_{i}$ with the current question $Q^{n}_{i}$, replacing the reference to the bridging entity $A^{n}_{i}$ in $q^{n+1}_{i}$ with the current question $Q^{n}_{i}$, resulting in the complete $(n+1)$ hop question $Q^{n+1}_{i}$.
  5. 5.
    Answer Normalization: Categorize the answer $A^{n+1}_{i}$ into three types: numerical values, time-related entities (e.g., years, dates), and strings. Construct standardized answer sets for each category.
Through this iterative process, we systematically develop multi-level reasoning question sets, ranging from two-hop to three-hop.
Comparison Dataset
This question type evaluates cross-image reasoning, as exemplified in Figure 2 (b). Generating these questions requires identifying entities across multiple images and utilizing external knowledge to formulate a comparative query. The construction process is as follows:
  1. 1.
    Entity Collection: Collect a large number of visual entities from the Wikipedia knowledge base. Use the embedding model NV-Embed-v2 (Lee et al. 2024) to perform semantic similarity matching based on the entity names and summary, selecting entities with high relevance. Use LLMs to perform semantic deduplication and remove pairs of entities that refer to the same concept.
  2. 2.
    Paragraph Matching: For the selected similar entity pairs, use a rule-based method to extract paragraphs with the same title that contain quantifiable numerical information, providing consistent background knowledge for subsequent comparative analysis.
  3. 3.
    Question Generation: Based on the entity pairs and their background knowledge paragraphs, prompt the LLMs to focus on quantifiable attributes and generate questions that compare the attributes of the two entities.
  4. 4.
    Quality Control: Perform automated validation to ensure that the questions are clearly stated and the answers are quantifiable and verifiable.
  5. 5.
    Multimodal Transformation: Replace the entity names in the questions with corresponding images to construct a multimodal reasoning scenario that forces the model to reason based on visual content.
  6. 6.
    Answer Normalization: Standardize numerical answers, ensuring that clear units are included in the question to support accurate evaluation.
Figure 3: Overview of the training pipeline for MMhops-R1.
MMhops Analysis
MMhops is the first large-scale dataset designed for multimodal multi-hop reasoning. As detailed in Table 2, the dataset comprises 31,117 samples, which include 20,483 unique questions, 8,832 distinct entities, and 28,256 images. A key feature of MMhops is its focus on complex reasoning chains; all samples require more than two reasoning hops that span both visual and textual modalities. Specifically, 70.8% of samples require three reasoning steps and 29.2% require four. Furthermore, all samples necessitate the integration of external knowledge. Linguistically, the average question length is 17.3 words, with concise answers averaging 1.6 words. Most answers are numerical, facilitating precise evaluation of the model’s reasoning ability.
As detailed in Table 1, existing Knowledge-based VQA (KVQA) datasets are largely confined to shallow reasoning, typically involving a single visual step and 1–2 textual reasoning hops. Consequently, they are insufficient for evaluating complex, multi-step reasoning abilities. MMhops dataset incorporates multi-image inputs, which necessitates 1–2 steps of cross-image relational reasoning. Furthermore, we extend the textual reasoning depth to 2–3 hops via a scalable, iterative pipeline. Collectively, these enhancements result in a total reasoning depth of 3–4 steps, establishing MMhops as a more challenging and practical benchmark to drive progress in advanced multimodal reasoning.
Dataset Splits
We split the MMhops dataset into training, validation, and test sets with a 7:1:27{:}1{:}2 ratio using stratified sampling based on reasoning depth and question type.
Methodology
Problem Formulation
We consider the task of answering a question $Q$ based on a collection of images $\mathcal{I}=\{I_{1},\ldots,I_{n}\}$. The policy model $\pi_{\theta}$ can leverage a set of external retrievers $\mathcal{R}=\{R_{I},R_{T}\}$. $R_{I}$ is an image retriever that, given a query image, returns the information about the most similar image. $R_{T}$ is a text retriever that, given a text query, returns the top-$k$ most relevant passages. The model’s action space is defined as $\mathcal{A}=\{a_{t},a_{is},a_{ts},a_{a}\}$, where $a_{t}$ represents thinking and reasoning based on all current inputs, $a_{is}$ and $a_{ts}$ invoke the retrievers $R_{I}$ and $R_{T}$, respectively, and $a_{a}$ terminates the process by generating the final response. The training pipeline for our policy model is illustrated in Figure 3.
Rollout with Multi-Step Retrieval
At each time step $t$, the policy model $\pi_{\theta}$ first performs thinking based on the current state $s_{t}$ (containing historical interaction information), then selects the next action from the action set $\{a_{is},a_{ts},a_{a}\}$. The specific action execution mechanisms are as follows:
  • Image Retrieval ($a_{is}$): The policy issues a query to an image retriever $R_{I}$ by generating target image indices, formatted within <image_search> and </image_search> tags. The retriever returns the corresponding image information as observation $o_{t+1}$.
  • Text Retrieval ($a_{ts}$): The policy generates a textual query, formatted within <text_search> and </text_search> tags, for a text retriever $R_{T}$, which returns the top-$k$ relevant passages as observation $o_{t+1}$.
  • Answer ($a_{a}$): The policy generates the final answer, formatted within <answer> and </answer> tags, based on the information gathered throughout the trajectory. This is a terminal action that concludes the episode.
If the policy generates an action with a malformed syntax or one outside $\mathcal{A}$, the environment provides a fixed penalty signal as the observation $o_{t+1}$ to encourage valid action generation. The rollout process terminates when the policy executes the answer action $a_{a}$ or a maximum of $T$ steps is reached. This interaction generates a trajectory $\tau$, defined as a sequence of states, actions, and observations:
$\tau=\{(s_{0},a_{0},o_{1}),(s_{1},a_{1},o_{2}),\ldots,(s_{T},a_{T})\}.$ (1)
Reward Modeling
To guide the model’s generation, we design a composite reward function for MMhops-R1. This function comprises three components designed to promote correctness, structural clarity, and effective tool use. For a given trajectory $\tau$, the total reward is a weighted sum of these components.
  1. 1.
    Outcome Reward ($R_{\text{outcome}}$). This binary reward evaluates the correctness of the final answer. It is defined as $R_{\text{outcome}}(\tau)=1$ if the model’s answer matches the ground truth, and 0 otherwise.
  2. 2.
    Format Reward ($R_{\text{format}}$). This binary reward encourages adherence to the previously defined structured format. A trajectory receives $R_{\text{format}}(\tau)=1$ if all generated thoughts and actions are correctly formatted with their respective tags, and 0 otherwise.
  3. 3.
    Action Reward ($R_{\text{action}}$). This rewards effective tool use. A key aspect of our design is that this reward is gated by the overall success of the trajectory. It is only granted if the model both produces the correct final answer and adheres to the required format. This encourages the model to learn tool-use policies that directly contribute to successful outcomes. The reward is defined as:
    $R_{\text{action}}(\tau)=R_{\text{outcome}}(\tau)\cdot R_{\text{format}}(\tau)\cdot R_{\text{tool}}(\tau)$ (2)
    where $R_{\text{tool}}(\tau)$ is a separate reward, defined as the number of syntactically correct tool invocations.
The total reward for a trajectory $\tau$ is a weighted sum of these components:
$R(\tau)=\alpha\cdot R_{\text{outcome}}(\tau)+\beta\cdot R_{\text{format}}(\tau)+\gamma\cdot R_{\text{action}}(\tau)$ (3)
where $\alpha,\beta,\text{and }\gamma$ are non-negative hyperparameters that balance the contribution of each component.
Bridging
Method Base Model Retriever String Numerical Time Overall Comparison
Closed-sourced model
GPT-4o-mini (Hurst et al. 2024) 29.67 20.54 26.08 23.80 7.05
GPT-4o (Hurst et al. 2024) 41.66 33.60 39.28 36.62 8.76
Gemini-2.5-flash (Comanici et al. 2025) 51.08 43.51 50.10 46.58 23.18
Gemini-2.5-pro (Comanici et al. 2025) 58.80 50.83 57.32 53.98 29.39
Direct Answer
Zero-shot Qwen2.5-vl-7B-Instruct 24.21 15.24 26.60 19.53 6.20
Zero-shot Qwen2.5-vl-72B-Instruct 37.51 32.11 37.32 34.39 7.59
Multi-hop RAG (Text-only)
Search-r1 (Jin et al. 2025) Qwen2.5-7b-Instruct Caption, Text 14.45 23.05 17.85 19.98 6.62
Self-Ask (Press et al. 2022) GPT-4o Caption, Text 27.59 31.41 31.13 30.42 18.27
Multimodal RAG
Vanilla mRAG Qwen2.5-vl-7B-Instruct Text 14.37 14.65 14.95 14.63 3.95
Vanilla mRAG Qwen2.5-vl-7B-Instruct Image, Text 26.52 25.68 28.97 26.49 9.72
EchoSight (Yan and Xie 2024) LLaMA3 Image, Text 19.14 11.83 11.86 13.63 4.81
OmniSearch (Li et al. 2024) GPT-4o Image, Text 31.02 49.77 36.5 42.65 17.02
MMhops-R1 (Ours) Qwen2.5-vl-7B-Instruct Image, Text 44.66 55.33 47.94 51.35 22.01
Table 3: Main results on MMhops.
Objective Function
To optimize our policy $\pi_{\theta}$ using a composite reward signal, we adapt the objective function from DAPO (Yu et al. 2025). This objective is coupled with a dynamic sampling strategy that filters generated response groups. Specifically, we enforce that each group of $G$ responses must contain at least one factually correct sample, as stipulated by the constraint in our objective. This design enables the policy to optimize for procedural correctness by creating non-zero advantages ($\hat{A}_{i,t}$) from process-based rewards, such as format adherence ($R_{\text{format}}$), even when the final outcome is already correct. Our full optimization objective is formulated as:
$\begin{split}J(\theta)&=\mathbb{E}_{(q,a)\sim\mathcal{D},\{o_{i}\}^{G}_{i=1}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q;R)}\\ &\quad\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\min\left(r_{i,t}(\theta)\hat{A}_{i,t},\right.\right.\\ &\quad\left.\left.\text{clip}\left(r_{i,t}(\theta),1-\epsilon_{\text{low}},1+\epsilon_{\text{high}}\right)\hat{A}_{i,t}\right)\right]\\ \text{s.t.}\quad&0<\left|\{o_{i}\mid\text{is\_equivalent}(a,o_{i})\}\right|.\end{split}$ (4)
where
$r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i},\leq t;R)}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i},\leq t;R)},\text{and}$
$\hat{A}_{i,t}=\frac{R_{i}-\text{mean}(\{R_{i}\}^{G}_{i=1})}{\text{std}(\{R_{i}\}^{G}_{i=1})}$
Where $R$ represents the retriever, and the model samples while interacting with multiple retrievers $R=\{R_{I},R_{T}\}$.
Loss Masking for External Observations
The observation $o_{t+1}$ at each timestep $t$ includes tokens from external sources, such as results from the image ($R_{I}$) and text ($R_{T}$) retrievers and environmental feedback on invalid actions. Since these tokens are not generated by the policy model, we mask them from the loss computation during policy optimization. This ensures that the optimization objective is confined to the model’s own generated reasoning and action tokens, improving training stability.
Experiments
Experimental Settings
Implementation Details
We optimize our policy using the Verl framework (Sheng et al. 2024), employing Qwen2.5-VL-7B-Instruct (Bai et al. 2025) as the backbone model. The model is trained for a single epoch on the MMhops dataset with a constant learning rate of $1\times 10^{-6}$. During policy optimization, we use a batch size of 256 and a group size of 8. Our knowledge base for retrieval comprises 100K Wikipedia articles, each accompanied by an image. For image retrieval, we utilize the CLIP-ViT-L/14@336px model (Radford et al. 2021). For text retrieval, we employ the E5 model (Wang et al. 2022) to fetch the top-3 most relevant passages for each query. The maximum number of interaction turns with the knowledge base is set to 4 during both training and inference. The hyperparameters $\alpha$, $\beta$, and $\gamma$ for the reward function are set to 1.0, 1.0, and 0.25, respectively.
Evaluation Metrics
We adopt the evaluation protocol from INFOSEEK (Chen et al. 2023), categorizing answers into three types: String, Time, and Numerical. For String answers, we report Exact Match (EM) accuracy. For Time answers, we employ EM with a tolerance of $\pm 1$ year. For Numerical answers, a prediction is deemed correct if it falls within a $\pm 0.1$ margin of the ground truth or achieves an Intersection-over-Union (IoU) of at least 50%. The overall score is the weighted average of the accuracies for each type. We report performance on the Test set, with a breakdown for bridging and comparison questions.
Comparison with SOTAs
To enable a comprehensive comparison with existing approaches, we evaluate four categories of models: advanced open-source general-purpose multimodal large models with direct answer generation, single-modal text-only multi-hop RAG methods (which convert images into descriptions and combine them with the question as input), multimodal RAG methods, and benchmark closed-source MLLMs. For fairness, all compared methods share the same image and text retriever as ours. Detailed results are presented in Table 3, with key findings summarized as follows:
1. Rich domain knowledge and strong reasoning capabilities are essential for solving multimodal multi-hop problems. However, general-purpose open-source MLLMs are relatively weak in both aspects, posing challenges for them to generalize to this task. Even the 72B Qwen2.5-VL model falls short by 16.96% and 14.42% in overall accuracy on bridging and comparison questions, respectively, compared to our 7B-based model.
2. Incorporating visual information is fundamental to effective multimodal reasoning. Text-only multi-hop RAG methods are unable to access critical visual information, making it difficult to perform appropriate knowledge retrieval and accurate reasoning for multimodal multi-hop problems. Specifically, the state-of-the-art method Self-Ask, which significantly boosts base model performance (e.g., GPT-4o) on textual multi-hop tasks, even shows an overall performance drop on comparison questions in MMhops compared to GPT-4o alone (30.42% vs. 36.62%).
3. Accurate multi-turn reasoning and retrieval interactions are critical to successfully solving multimodal multi-hop problems. Existing multimodal RAG methods, such as those designed for KB-VQA, are tailored to single-hop tasks and lack the ability to properly decompose multimodal multi-hop questions into sequential reasoning and retrieval steps, thereby limiting their answer accuracy. Even with the support of GPT-4o’s OmniSearch, the overall accuracy on bridging and comparison questions remains 9.7% and 4.99% lower than ours.
4. Closed-sourced commercial MLLMs remain the performance ceiling but still fall short of real-world applicability. Gemini-2.5-Pro, which has likely undergone reasoning-specific optimization and large-scale pretraining, outperforms our method but answers only about half of the bridging questions correctly, with lower accuracy on comparison questions. This underscores that multimodal multi-hop RAG remains largely unexplored.
Model INFOSEEK
Unseen Q Unseen E Overall
CLIP-PaLM (Chen et al. 2023) 22.7 18.5 20.4
CLIP-FiD (Chen et al. 2023) 23.3 19.1 20.9
Wiki-LLaVA (Caffagni et al. 2024) 30.1 27.8 28.9
EchoSight (Yan and Xie 2024) 31.3
MMhops-R1 33.8 32.6 33.2
Table 4: Comparison on INFOSEEK. Q: Question, E: Entity.
Model PaLI PaLM GPT-3 MMhops-R1
Two hop 14.7 22.8 18.7 23.3
Table 5: Comparison on E-VQA.
Cross-dataset Generalization Verification
To verify the generalizability of the proposed method, we evaluate it on two widely used knowledge-based VQA datasets: INFOSEEK (Chen et al. 2023) and E-VQA (Mensink et al. 2023). Results on INFOSEEK show the effectiveness of MMhops-R1 on multimodal single-hop questions, while its performance on two-hop questions in E-VQA confirms its generalization ability to multi-hop reasoning.
Ablation Studies
Bridging
Method String Numerical Time Overall Comparison
MMhops-R1 44.66 55.33 47.94 51.35 22.01
w/o $R_{\text{action}}$ 39.74 51.19 46.8 47.57 20.62
w/o $R_{\text{format}}$ 43.12 53.64 47.73 49.97 14.42
w/o $R_{\text{format}}$$R_{\text{action}}$ 40.43 42.68 40.62 41.75 13.03
Table 6: Effect of $R_{\text{outcome}}$, $R_{\text{format}}$ and $R_{\text{action}}$.
Bridging
Method String Numerical Time Overall Comparison
5 44.20 56.13 49.18 51.92 20.09
4 44.66 55.33 47.94 51.35 22.01
3 40.05 51.79 46.70 47.97 13.78
2 30.75 46.98 30.13 39.93 9.83
Table 7: Effect of the maximum retriever interaction count.
Effect of $R_{\text{outcome}}$$R_{\text{format}}$ and $R_{\text{action}}$.
As shown in Table 6: (1) Removing either the retrieval reward $R_{\text{action}}$ or the format reward $R_{\text{format}}$ leads to a notable performance drop, particularly on comparison questions; (2) Removing both $R_{\text{action}}$ and $R_{\text{format}}$ results in an even greater decline. These findings indicate that encouraging appropriate retrieval, enforcing correct feedback formats, and imposing strong constraints on answer precision all contribute positively to model performance.
Effect of Number of Interaction Rounds.
To demonstrate that the MMhops dataset indeed requires multi-step reasoning and RAG interaction for problem solving, we report model performance in Table 7 under maximum rounds constrained to 2, 3, 4, and 5. As the number of rounds increases from 2 to 4, overall performance consistently improves, while further increasing to 5 yields no significant gains but introduces more computational overhead. Therefore, four-step reasoning is most suitable for MMhops.
Conclusion
In this work, we introduce the first large-scale multimodal multi-hop reasoning dataset MMhops to evaluate models’ capabilities in multi-turn interactive reasoning and external knowledge utilization, and extensive experiments show that existing MLLMs struggle on MMhops. To address this, we further propose a novel reinforcement learning-based framework MMhops-R1 for multimodal reasoning and RAG interaction. Results demonstrate that MMhops-R1 substantially outperforms existing methods by effectively leveraging reasoning and retrieval capabilities. The code, dataset, and model weights will be open-sourced to encourage future research on the multimodal multi-hop reasoning task.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. U24A20331, No. 62302501), the Beijing Natural Science Foundation (No. L251005, No. L243015) and the Key Research and Development Program of Xinjiang Uyghur Autonomous Region (No. 2023B01005).
Supplementary Materials
This section provides the detailed prompts used to construct the MMHops dataset. The following prompts were carefully engineered for state-of-the-art large language models like GPT-4o. They incorporate specific constraints and few-shot examples to ensure high-quality, structured output for both Bridging and Comparison question types.
Bridging Question
Prompt for Sub-Question Generation
Bridging questions are generated by creating a new sub-question based on a bridging entity. The following prompt template was used for this task.
Prompt: Bridging Sub-question Generation Goal Given a Wikipedia entity and its corresponding Wikipedia content, you need to formulate a question about the entity or its attributes, provide the answer, and indicate the source sentence for the answer. Limitations Question: 1. The question must be about the given entity, inquiring about its attributes, such as the birth date of a person or the land area of a country. 2. The original name of the given entity must appear in the question and cannot be a variant. It should only appear once; if it needs to appear multiple times, pronouns should be used. 3. The question must not introduce additional information that allows the answer to be inferred from perspectives other than the given entity, including: (1) When asking about the given entity, do not focus on another independent entity in such a way that the question could be answered by only considering the other entity. (2) Avoid introducing highly restrictive or directly answerable expressions. 4. The question must unambiguously point to a unique answer. If the question is time-sensitive, a time frame should be included. Answer: 1. The answer should not be open-ended; it should be a specific noun, numerical value, or exact date, expressed in one word or a short phrase. 2. Avoid questions that lead to incomplete answers or have multiple valid answers. 3. The answer must be unambiguously derivable from the evidence sentence. Evidence Sentence: 1. The evidence sentence must form a complete reasoning chain. Both the subject of the question and the answer must appear in the evidence sentence.   Examples Example 1: Wikipedia entity: Bean Wikipedia content: Taxonomy.The Fabaceae are placed in the order Fabales… The family now includes six subfamilies:Cercidoideae: 12 genera and  335 species… Output: 1. Question: How many subfamilies are included in the Fabaceae family? 2. Answer: Six 3. Evidence sentence: The family now includes six subfamilies:Cercidoideae: 12 genera and  335 species. Example 2: Wikipedia entity: Russia Wikipedia content: Culture.Russian culture has been formed by the nation’s history… Russia is home to 30 UNESCO World Heritage Sites, 19 out of which are cultural… Output: 1. Question: How many UNESCO World Heritage Sites are there in Russia? 2. Answer: 30 3. Evidence sentence: Russia is home to 30 UNESCO World Heritage Sites, 19 out of which are cultural; while 27 more sites lie on the tentative list.   Real Data Wikipedia entity: {ENTITY_ NAME} Wikipedia content: {WIKIPEDIA_CONTENT} Output Format You need to strictly follow the format below for output, and do not output anything other than the specified content. Only generate one question and answer pair: 1. Question: [Your generated question] 2. Answer: [Your generated answer] 3. Evidence sentence: [The evidence sentence]
Prompt for Question Quality Control
After generating a sub-question, a quality control step is performed to ensure it does not contain information leakage and adheres to structural requirements. The following prompt was designed for this binary classification task.
Prompt: Question Quality Control Goal Given a Wikipedia entity and a question about this entity, determine whether the question meets the following requirements. If it does, reply with True; if it does not, reply with False. Requirements 1. The question must focus on the given named entity itself and its attributes, not on derivatives of the entity. The entity name must appear in full and uninterrupted in the question. 2. The question should not reveal additional information, such as restrictive descriptions, which would allow a model to bypass the given entity and answer the question. 3. The question should not introduce independent new entities that allow it to be answered without using the given entity. Guiding Principle: The question must be unanswerable if the given entity name is removed from it.  True Examples Example 1: Entity name: Fabaceae Question: What type of fruits do Fabaceae produce after fertilization? Output: True Example 2: Entity name: Spain Question: How many World Heritage Sites does Spain have? Output: True Example 3: Entity name: Robert Adler Question: At what age did Robert Adler die? Output: True   False Examples Example 1: Entity name: AM General Question: In what year did AM General begin preliminary design work on the M998 Series High Mobility Multi-Purpose Wheeled Vehicle? Explain: The introduction of the ”M998 Series…” allows one to obtain the answer by focusing only on this part, making it possible to skip the given entity ”AM General.” Output: False Example 2: Entity name: Inuit Question: What is the name of the northernmost city in the United States where the Inuit of Alaska live? Explain: The description ”the northernmost city in the United States” allows the question to be answered directly without focusing on the entity ”Inuit.” Output: False Example 3: Entity name: Apple Inc. Question: When did Steve Jobs announce that Apple Computer, Inc. would be known as ”Apple Inc.”? Explain: ”Steve Jobs” and ”Apple Computer, Inc.” can be used to infer the answer without needing to focus on the given entity ”Apple Inc.” Output: False   Real Data Entity name: {ENTITY_NAME} Question: {QUESTION_TEXT} Output Format Please study the given examples and verify the provided real data. Output only True or False without any additional content.
Prompt for Question Merging
To construct complex, multi-hop questions from simpler components, a merging step is essential. This prompt is designed to take an initial question and a follow-up question about its answer (the bridging entity) and combine them into a single, coherent multi-hop question.
Prompt: Question Merging Goal Given an original question, its answer (which serves as a ”bridging entity”), and a second question about that entity, merge the two questions into a single, new question. Requirements 1. Substitution: Replace the bridging entity in the second question with a descriptive clause derived from the original question. The name of the bridging entity must not appear in the final output. 2. Cohesion: The output must be a single, grammatically complete question, not two separate clauses. 3. Fluency: The merged question must be grammatically correct and sound natural in English.   Examples Example 1: Original question: Who is the manufacturer of this vehicle? Answer/Bridge entity: AM General Second question: In what year did AM General begin producing purpose-built taxi-cabs? Two-hop question: In which year did the manufacturer of this vehicle begin producing purpose-built taxi-cabs? Example 2: Original question: What is this person’s place of birth? Answer/Bridge entity: Brooklyn Second question: When did the Brooklyn Museum open? Two-hop question: When did the museum in this person’s place of birth open? Example 3: Original question: Where is this bird native to? Answer/Bridge entity: New Zealand Second question: What percentage of New Zealand’s indigenous vascular plants are endemic? Two-hop question: What percentage of indigenous vascular plants in the area where this bird is native are endemic? Example 4: Original question: What country does this building belong to? Answer/Bridge entity: Cambodia Second question: What is the temperature range in Cambodia in degrees Celsius? Two-hop question: What is the temperature range in the country to which this building belongs in degrees Celsius?   Real Data Original question: {ORIGINAL_QUESTION} Answer/Bridge entity: {BRIDGE_ENTITY} Second question: {SECOND_QUESTION} Output Format Please study the given examples and merge the two questions according to the requirements. Output only the final merged question, with no additional text or labels.
[Your generated multi-hop question]
Comparison Question
Prompt for Comparison Question Generation
This prompt is designed to generate high-quality comparative questions from two distinct text sources. It focuses on creating questions that require a precise, data-driven comparison between two entities, ensuring the output is structured, verifiable, and directly supported by the provided evidence.
Prompt: Comparison Question Generation Goal Given two Wikipedia entities and paragraphs about these two entities, formulate a question comparing the two entities based on the content of the given paragraphs. Provide the answer and the sentence from which the evidence for the answer is derived. Requirements 1. The question must focus on comparing the same attribute of two given entities. It should not consider only a single entity, nor shift the comparison to other entities. The nature of the question must be comparative and not of any other type, such as summarization. 2. The question should focus on comparing two entities with respect to specific attributes such as numerical values, dates, or years. 3. The question description should be as detailed as possible, fully including the complete names of the two given entities (no splitting or abbreviations), and clearly specifying the expected answer format and units. 4. The answer must be a single word or phrase, not a list of attributes for the two entities. Numerical answers should be expressed using Arabic numerals with units. The supporting evidence must be a complete sentence taken directly from the original text. 5. If a valid question cannot be generated based on the given content, return False directly.   Examples Example 1: Successful Numerical Comparison Entity 1: Gull-billed tern Description: This is a fairly large and powerful tern… It is 33 - 42 (cm) in length and 76 - 91 (cm) in wingspan. Body mass ranges from 150 - 292 (g)… Entity 2: Caspian tern Description: It is the world’s largest tern with a length of 48–60 (cm), a wingspan of 127–145 (cm) and a weight of 530 - 782 (g)… Output: 1. Question: What is the difference in centimeters between the maximum wingspans of the Gull-billed Tern and the Caspian Tern? 2. Answer: 54 cm 3. Evidence sentences: Entity 1: It is 33 - 42 (cm) in length and 76 - 91 (cm) in wingspan. Entity 2: It is the world’s largest tern with a length of 48 – 60 (cm), a wingspan of 127 – 145 (cm) and a weight of 530 - 782 (g). Example 2: Invalid Comparison Entity 1: Château de Termes History: …the castle only fell to Simon de Montfort after a siege lasting four months, from August to November 1210… Entity 2: Château de Saissac History: Based on historical texts, it can be dated to at least 960… Note: A valid comparison is not possible because the entities do not share a comparable attribute. One has a specific siege date, while the other has a date of origin. Output:
False
  Real Data Entity 1: {ENTITY_1_NAME} {ENTITY_1_DESCRIPTION} Entity 2: {ENTITY_2_NAME} {ENTITY_2_DESCRIPTION} Output Format Strictly follow the format from the examples. Do not output any other content. 1. Question: … 2. Answer: … 3. Evidence sentences:… or
False
Prompts for Comparison Question Verification
To ensure the quality of complex comparative questions, we employ a verification prompt that assesses them against strict logical and evidential criteria. This prompt validates the question’s structure, the answer’s format, and the verifiability of the claim against the provided evidence, as detailed below.
Prompt: Comparison Question Verification Goal Given two Wikipedia entities, a question-answer pair, and an evidence sentence, verify whether the question-answer pair meets the following requirements. Reply with True if it does, otherwise False. Requirements 1. Question Integrity: The question must compare the two given entities, with both entity names fully appearing in the question. The question type must only compare numerical-related information for the given entities. 2. Answer Format: The answer must be a numerical value with a unit, or a numerical range. 3. Verifiability: The question must be clearly stated. Based on the supporting sentences, the question and answer must be verifiable without ambiguity. The attributes being compared in the question must be explicitly found in the supporting sentences; vague attribute references in the question do not meet this requirement.   Examples Example 1: Entity 1: Blanding’s turtle Entity 2: Northern map turtle Question: Which turtle has a larger maximum clutch size, Blanding’s turtle or Northern map turtle, and what is that maximum size? Answer: Northern map turtle, 20 eggs Evidence Sentences: Blanding’s turtle: ”The clutch size varies from region to region. In New York, the clutch size ranges from 5–12 eggs with an average of eight.”
Northern map turtle: ”The size of the clutch is between six and 20.”
Output: False
Example 2: Entity 1: Glacier National Park (U.S.) Entity 2: Yellowstone National Park Question: Which national park has a greater number of documented mammal species, and by how many species? Answer: Glacier National Park (U.S.) by 2 species Evidence Sentences: Glacier National Park (U.S.): ”Sixty-two species of mammals have been documented…”
Yellowstone National Park: ”There are almost 60 species of mammals in the park…”
Output: False
Example 3: Entity 1: Humvee Entity 2: Hummer H1 Question: What is the difference in years between the start of production of the Humvee and the Hummer H1? Answer: 9 years Evidence Sentences: Humvee: ”AM General was awarded an initial contract in 1983 for 2,334 vehicles, the first batch of a five-year contract that would see 55,000 vehicles delivered to the U.S. military.”
Hummer H1: ”Originally released in the civilian market March 14, 1992, the Hummer H1 became known from photographs published during Operation Desert Storm and the enthusiastic promotion by actor Arnold Schwarzenegger.”
Output: False
The evidence is not sufficient and direct enough.
Example 4: Entity 1: Crescent City Connection Entity 2: Blue Water Bridge Question: What is the difference in years between the opening of the first spans of the Crescent City Connection and the Blue Water Bridge? Answer: 20 years Evidence Sentences: Crescent City Connection: ”Construction of the first span began in November 1954, and the bridge opened in April 1958 as the Greater New Orleans Bridge.”
Blue Water Bridge: ”The first bridge was fully opened to traffic on October 10, 1938.”
Output: True
Example 5: Entity 1: Rumex obtusifolius Entity 2: Rumex crispus Question: What is the difference in centimeters between the maximum heights of Rumex obtusifolius and Rumex crispus? Answer: 50 cm Evidence Sentences: Rumex obtusifolius: ”Rumex obtusifolius is a perennial herbaceous flowering plant that grows to a height of 40 to 150 (cm).” Rumex crispus: ”The plant produces an inflorescence or flower stalk that grows to 1.5 (m) high.” Output: False The answer contains factual errors and inaccurate calculation results.   Real Data Entity 1: {ENTITY_1} Entity 2: {ENTITY_2} Question: {QUESTION} Answer: {ANSWER} Evidence Sentences: {EVIDENCE_SENTENCE} Output Format Strictly follow the format from the examples. The output must be the single word True or False.
Prompts for Comparison Question Rewriting
To adapt our comparative questions for multimodal contexts, we use a prompt to rewrite them into a format that references images instead of named entities. This process instructs the LLM to replace entity names with generic references like ”the building in the first image” while preserving the original question’s comparative intent, as detailed below.
Prompt: Comparison Question Rewriting Goal Given a question that compares two entities, your task is to rewrite the question to eliminate the entity names, replacing them with references to two images, while keeping the question’s meaning unchanged. Your response should be the rewritten question. Requirements 1. Preserve Meaning: The rewritten question must not include the original entity names, but its meaning must remain the same and still point to the given answer. To ensure the answer format remains as entity names, add ”(answer the name)” where appropriate. 2. Image Referencing: The two entities correspond to two images: entity 1 is in the first image, and entity 2 is in the second. Replace references to the entities in the question with references to the images, either individually or collectively. 3. Use Entity Type: In the rewritten question, you can use the type of the entity (e.g., car, airplane, animal, plant, building) to replace the entity name for clarity. 4. Clarity and Brevity: The rewritten question should be concise and natural, avoiding complex or awkward phrasing.   Examples Example 1: Entity 1: Common toad Entity 2: Common frog Question: How many more eggs can a Common Toad lay compared to a Common Frog at maximum capacity in eggs? Answer: 30,000 eggs Output: How many more eggs can the animal in the first image lay compared to the second image at maximum capacity in eggs? Example 2: Entity 1: Trumpetfish Entity 2: Cornetfish Question: What is the difference in meters between the maximum lengths of the Trumpetfish and the Cornetfish? Answer: 1 m Output: What is the difference in meters between the maximum lengths of the fish in the first image and the second image?   Real Data Entity 1: {ENTITY_1} Entity 2: {ENTITY_2} Question: {QUESTION} Answer: {ANSWER} Output Format The output should be only the rewritten question.
Answer Processing and Normalization
Prompt for Answer Verification and Classification
To ensure the generated question-answer pairs are valid and to categorize them for analysis, a verification and classification prompt is used. This step confirms that the answer is derivable from the evidence and assigns it a specific data type.
Prompt: Answer Verification and Classification Goal Given a question-answer pair and the corresponding evidence sentence, verify whether the pair meets the requirements and classify the type of the answer. Requirements 1. The answer can be precisely inferred from the evidence sentence. 2. The answer must be a specific number, numerical range, year, date, or string. It cannot be a vague concept. Classification Based on the answer, classify it into one of the following three types. If it does not meet the requirements or is a vague concept, classify it as False. 1. For date-type answers, respond with Time. This is limited to specific years (e.g., 1897) and dates (e.g., 1 January 1981). General ranges like ”the 16th century” are classified as String. 2. For numerical-type answers, respond with Numerical, which includes specific numbers or specific numerical ranges. Answers of this type generally include a unit of measurement; if no unit is provided, the answer should represent a count or percentage. 3. For string-type answers, respond with String. This applies to any other specific, non-numerical, non-date answer (e.g., Zebedee, 4th, 16th century). For answers that do not meet the above requirements, cannot be inferred from the given evidence sentence, or are vague concepts or ranges, respond with False.  Time Examples Example 1: Question: When did the Brooklyn Museum open? Answer: 1897 Evidence Sentences: The Brooklyn Museum, opened in 1897, is New York City’s second-largest public art museum. Output: Time   Numerical Examples Example 1: Question: What percentage of all genera of land plants were angiosperms in the Maastrichtian? Answer: 50% to 80% Evidence Sentences: From 50% to 80% of all genera of land plants were angiosperms… Output: Numerical Example 2: Question: How many living subdivisions does Cryptobranchoidea have? Answer: Two Evidence Sentences: It has two living subdivisions, Cryptobranchidae… and Hynobiidae… Output: Numerical   String Examples Example 1: Question: Who was John the Apostle’s father? Answer: Zebedee Evidence Sentences: John the Apostle was the son of Zebedee and the younger brother of James the Great. Output: String Example 2: Question: When did the dominance of Gothic architecture begin to wane? Answer: 16th century Evidence Sentences: Beginning in the 16th century… the dominance of Gothic architecture began to wane. Output: String   False Examples Example 1: Question: How many members of parliament are elected in Turkey? Answer: Five to six hundred thousand Evidence Sentences: Out of a population of 9.5 million, it is estimated that five to six hundred thousand people sing in choirs. Explain: The evidence sentence does not support the question, and the answer ”five to six hundred thousand” is presented as a vague estimate, not a precise range. Output: False   Real Data Question: {QUESTION} Answer: {ANSWER} Evidence Sentences: {EVIDENCE_SENTENCE} Output Format Please classify the given real data according to the instructions. Output only one of the specified classification types: Time, Numerical, String, or False.
Prompt for Numerical Answer Normalization
To create a robust and machine-readable evaluation set for questions with numerical answers, the raw text answers must be standardized. This prompt automates the process of validating, extracting, and normalizing numerical data, and refines the corresponding question to ensure clarity for automated evaluation.
Prompt: Numerical Answer Normalization Goal Given a question-answer pair and its evidence sentence, standardize the answer into a machine-readable format. This involves extracting the numerical value and its unit, and rewriting the question to explicitly request that unit. Processing Steps 1. Extract & Standardize Value: Extract the numerical value(s), converting words (e.g., ”Seven”) to digits (e.g., ”7”). Remove all units and formatting (e.g., commas, symbols). For a range, list the lower and upper bounds. The result should be a list of numbers. 2. Identify Unit: Extract the unit of measurement (e.g., %, million, °C). If the answer is a simple count, the unit is None. 3. Rewrite Question: Modify the original question, if necessary, to ensure it explicitly asks for the answer in the identified unit (e.g., ”What is the population?” becomes ”What is the population in millions?”).   Examples Example 1: Question: What percentage of Germany’s health care system was government-funded according to the World Health Organization? Answer: 77% Evidence Sentences: According to the World Health Organization (WHO), Germany’s health care system was 77% government-funded and 23% privately funded. Output: 1. Value: [77] 2. Unit: 3. Question: What percentage of Germany’s health care system was government-funded according to the World Health Organization? Example 2: Question: What is the population of Denmark as of 2022? Answer: 5.91 million Evidence Sentences: As of 2022, it had a population of 5.91 million (1 August 2022)… Output: 1. Value: [5.91] 2. Unit: million 3. Question: What is the population of Denmark in millions as of 2022? Example 3: Question: What is the temperature range in Cambodia? Answer: 21 to 35 (°C) Evidence Sentences: Cambodia has a temperature range from 21 to 35 (°C) and experiences tropical monsoons. Output: 1. Value: [21, 35] 2. Unit: °C 3. Question: What is the temperature range in Cambodia in degrees Celsius?   Real Data Question: {QUESTION} Answer: {ANSWER} Evidence Sentences: {EVIDENCE_SENTENCE} Output Format Strictly follow the format from the examples. The output must be either the three-line structure or the single word False. Do not output any other content. 1. Value: […] 2. Unit: … 3. Question: … or False
Prompt for String Answer Normalization
To robustly evaluate answers that are strings (e.g., names, places, concepts), it is necessary to account for all valid, synonymous expressions present in the source text. This prompt standardizes string answers by generating a comprehensive list of all equivalent terms found in the evidence.
Prompt: String Answer Normalization Goal Given a question-answer pair and the corresponding evidence sentence, standardize the answer, write out all synonymous expressions of the answer, and generate an answer list for evaluation. Processing Steps 1. Validate Answer: First, verify that the answer is a string and can be fully and accurately derived from the provided evidence sentence. If either condition is not met, the output must be False. 2. Generate Synonym List: If the answer is valid, identify and extract all equivalent expressions, names, or abbreviations for the answer from the evidence. Collate these into a single list of strings. All strings in the list should be in lowercase. If no synonyms are found, the list should contain only the original answer in lowercase.   Real Data Question: {QUESTION} Answer: {ANSWER} Evidence Sentences: {EVIDENCE_SENTENCE} Output Format Strictly follow the format from the examples. The output must be either the Answer_eval list or the single word False. Do not output any other content. Answer_eval: […] or False
Prompts for Time-based Answer Normalization
For time-based answers, this process involves standardizing various formats into a canonical form and generating an extensive list of acceptable variants for flexible matching, as detailed in the prompt below.
Prompt: Time-based Answer Normalization Goal Given a question-answer pair and the corresponding evidence sentence, standardize the answer and generate an evaluation list for the answer. Create the following two contents: Answer and Answer_eval. Processing Steps 1. Check if the answer is a specific year or a specific date. If it is, proceed to the next steps; if not, return False. 2. Verify if the answer can be inferred from the evidence sentence. If it can, fill the ”answer” list with the given answer, including only the year or date without any units. 3. Generate Answer_eval: If the answer is a specific year, include the exact year in the ”Answer_eval” list and add a range of plus and minus one year to create an evaluation list. If the answer is a specific date that includes a year, first include all standard expressions of the date in the ”Answer_eval” list. Then, adjust the year by plus and minus one and add all possible expressions for these adjusted dates. Finally, remove the date and list all possible years. If the answer is a specific date that does not include a year, list all possible standard expressions of the date in the ”Answer_eval” list. 4. If the above steps cannot be completed, return false directly.   Examples Example 1: Question: When did the Brooklyn Museum open? Answer: 1897 Evidence Sentences: The Brooklyn Museum, opened in 1897, is New York City’s second-largest public art museum. Output: 1. Answer: ["1897"] 2. Answer_eval: ["1897", "1896", "1898"] Example 2: Question: When did Greece become the tenth member of the European Communities? Answer: 1 January 1981 Evidence Sentences: Greece became the tenth member of the European Communities (subsequently subsumed by the European Union) on 1 January 1981, ushering in a period of sustained growth. Output: 1. Answer: ["1 January 1981"] 2. Answer_eval: ["1 January 1981", "January 1 1981", "1981 January 1", "1 1 1981", "1981 1 1", "1981", "1 January 1980", "January 1 1980", "1980 January 1", "1 1 1980", "1980 1 1", "1980", "1 January 1982", "January 1 1982", "1982 January 1", "1 1 1982", "1982 1 1", "1982"] Example 3: Question: When is Russia’s Unity Day celebrated? Answer: 4 November Evidence Sentences: Unity Day on 4 November, commemorating the 1612 uprising which marked the end of the Polish occupation of Moscow. Output: 1. Answer: ["4 November"] 2. Answer_eval: ["4 November", "November 4", "4th November", "November 4th", "4 11", "11 4"] Example 4: Question: When did Denmark stop issuing new licences for oil and gas extraction? Answer: December 2020 Evidence Sentences: Denmark stopped issuing new licences for oil and gas extraction in December 2020. Output: 1. Answer: ["December 2020"] 2. Answer_eval: ["December 2020", "2020 December", "12 2020", "2020 12", "2020", "December 2021", "2021 December", "12 2021", "2021 12", "2021", "December 2019", "2019 December", "12 2019", "2019 12", "2019"]   Real Data Question: {QUESTION} Answer: {ANSWER} Evidence Sentences: {EVIDENCE_SENTENCE} Output Format Strictly follow the format from the examples. The output must be either the two-line structure or the single word False. Do not output any other content. 1. Answer: [...] 2. Answer_eval: [...] or False